# Blind and Topological Interference Managements for Bistatic Integrated Sensing and Communication

### v4

@@ Abstract
Integrated sensing and communication (ISAC) systems provide significant enhancements in performance and resource efficiency compared to individual sensing and communication systems, primarily attributed to the collaborative use of wireless resources, radio waveforms, and hardware platforms. This paper focuses on the bistatic ISAC systems with dispersed multi-receivers and one sensor. Compared to a monostatic ISAC system, the main challenge in the bistatic setting is that the information messages are unknown to the sensor and therefore they are seen as interference, while the channel between the transmitters and the sensor is unknown to the transmitters. In order to mitigate the interference at the sensor while maximizing the communication degree of freedom, we introduce two strategies, namely, blind interference alignment and topological interference management. Although well-known in the context of Gaussian interference channels, these strategies are novel in the context of bistttic ISAC. For the bistatic ISAC models with heterogeneous coherence times or with heterogeneous connectivity, the achieved ISAC tradeoff points in terms of communication and sensing degrees of freedom are characterized. In particular, we show that the new tradeoff outperforms the time-sharing between the sensing-only and the communication-only schemes. Simulation results demonstrate that the proposed schemes significantly improve the channel estimation error for the sensing task compared to treating interference as noise at the sensor.
@@ I Introduction
Figure 1: The considered bistatic ISAC systems.
@@
Despite the fundamental differences in information processing principles between sensing and communication, significant research efforts have been devoted to exploring the essence of “integration” gains in ISAC systems. These studies focus on the information-theoretic trade-offs between communication and sensing performance, aiming to establish a theoretical foundation for their co-design. The information-theoretic framework for ISAC was first introduced in [3, 4], where the authors investigated a monostatic ISAC system with state-dependent discrete memoryless channels and an independent and identically distributed (i.i.d.) state sequence. The capacity-distortion region was characterized by maximizing conditional mutual information under specific constraints. To explicitly derive the closed-form on the trade-off between communication and sensing performances, Xiong et al. [5] analyzed the MIMO Gaussian channel and explored the Cramér-Rao Bound (CRB)-rate region, which identified two extreme trade-off points: the communication-optimal point and the sensing-optimal point, which further revealed two critical aspects of ISAC trade-offs: subspace trade-offs and deterministic-random trade-offs. For scenarios with fixed sensing states, where the state remains constant over a transmission block, the optimal trade-off between communication rate and state detection error exponent was characterized in [6]. Monostatic ISAC over interference channels was considered in [7, 8], where two interfering ISAC transmitters communicate with their own users or a common user while sensing estimation through received echo signals, and an improved random coding than the superposition coding for interference channels was proposed. Readers can refer to the surveys in [9, 10] for more details on the information theoretic progress for ISAC.
@@
The aforementioned researches focus on monostatic ISAC systems, where states are estimated by leveraging communication messages as side information. In contrast to monostatic sensing, bistatic sensing employs separated sensor and transmitter, which can avoid direct reflection or interference signals while covering larger areas. This approach exhibits advantages in the detection of low-altitude and ground targets, as well as in target detection utilizing multipath signals [11]. From an information-theoretic perspective [^1] The authors are fully aware of the fact that a whole host of problems related to the synchronization (time, frequency, phase) of transmitters and sensors play an important role in bistatic sensing. Nevertheless, these problems can be circumvented by appropriate system design. For example, we may imagine that base stations and sensor are part of the same network running a high precision synchronization protocol, which effectively eliminate the problem to the required degree of accuracy. In contrast, conveying the information bits sent to the users by multiple base stations without a suitable delay is a much harder task, since these messages depend on local (at the base station) scheduling decisions and the throughput of such traffic would be huge. Hence, it is reasonable to assume that the information messages are not know at the sensor, while the synchronization issues have been reduced to a degree that they can be neglected for the sake of analytical tractability. , research on bistatic ISAC mainly concentrates on two scenarios depending on the relative positions of communication receivers and sensors.
@@
  • For the bistatic ISAC model with dispersed receivers and sensor, a key challenge arises from the fact that communication messages are not necessary to be decoded by the sensor and thus act as its interference. [^2] In this paper, we mainly focus on how to cancel the interference from the communication message to the sensor. Thus another existing bistatic ISAC model with dispersed receivers and sensor, where the sensor knows the communication message (such as in [16]), is out of scope. The authors in [17] developed three achievable decoding-and-estimation strategies for sensors, blind estimation (e.g., by treating interference as noise as in [18]), by partial-decoding-based estimation (e.g., by successive interference cancellation (SIC) as in [19]), and full-decoding-based estimation (i.e., fully decoding the communication message and then estimating). Physical secrecy issue was considered in [20], where the sensor cannot obtain any information about the communication message while estimating the state.
@@
This paper considers the bistatic ISAC model with multiple transmitters, dispersed multi-receivers, and one sensor. In order to characterize the fundamental tradeoff between communication and sensing performance, we consider "homogeneous" metrics, communication degree of freedom (cDoF) and sensing degree of freedom (sDoF), representing the average number of effective transmissions and observations in each time slot, respectively. Two new interference management strategies are first introduced into the bistatic ISAC model, in order to perfectly eliminate interference to the sensor.333 The interference alignment strategy was introduced into ISAC systems in [21], by using the original interference alignment scheme in [22] with the assumption that the channels from the communication transmitters and the (transformed) channels from the radar transmitter to the sensor are all known. In our paper, we assume that any channel to the sensor is unknown.
@@
  • One uses the blind interference alignment (BIA) strategy originally proposed in [23]. We assume that each transmitter is connected to all the receivers and sensor through wireless channels, and the channel between the transmitter and the sensor is slow-fading and unknown to the transmitters, while the channels between the transmitter and communication receivers are fast-fading and known to the transmitter and receivers.444 The practical motivation is as follows: (i) Communication channels vary relatively quickly due to the mobility of communication users (with a coherence time of approximately 1-10 ms). Additionally, communication channels are periodically estimated via uplink/downlink (UL/DL) reciprocity and uplink pilots. Through these pilots, the transmitter obtains channel state information (CSI). For simplicity, we assume perfect CSI and do not consider pilot overhead or the actual channel estimation process. (ii) In contrast, sensing channels vary slowly because the sensors are stationary over few time slots. Furthermore, sensors and targets may be passive, meaning they do not transmit uplink pilots or provide channel state feedback. Therefore, even though the channel varies slowly, this CSI remains unknown to the transmitter. Note that the assumption that the CSI of communication channels is known and of sensing channels is unknown to the transmitters, has also been used in many information theoretic ISAC models such as [5, 17]. We consider three types of wireless channels, including interference channel, multi-user MISO (MU-MISO) channel, and multi-user MIMO (MU-MIMO) channel.
@@
  • The other strategy is the topological interference management (TIM) originally proposed in [24]. We follow the topological interference network model in [24, 25, 26], assuming that each transmitter is connected to a subset of the sensor and receivers through slow-fading wireless channels, while the transmitters only know the prior information of the network topology, instead of the CSI. In this model, cooperative and non-cooperative transmissions among the transmitters are respectively considered, depending on whether the transmitters have all communication messages or their own messages.
@@
By using the two interference management strategies, BIA and TIM, we propose new tradeoff points (in closed-form) between cDoF and sDoF, which are strictly better than the time-sharing between the two extreme sensing-optimal and communication-optimal points. The common ingredient of using the above two strategies into ISAC is to leverage the channel heterogeneity among the sensor and receivers to align interference at the sensor without knowing the CSI of the sensor channel, where the heterogeneity is on the coherence times for the first model, and on the connection topology for the second model. Finally, we perform simulations on practical ISAC system with QPSK modulation and additive white Gaussian noise. Considering the channel estimation error as the practical sensing metric, the proposed scheme improves the method by treating interference as noise at the sensor.
@@ II-A Channel Models
Consider a bistatic ISAC system, where ${\mathsf{K}}_{T}$ transmitters are equipped with ${\mathsf{m}}_{i}$ antennas where $i\in\mathcal{T}=\{1,\dots,{\mathsf{K}}_{T}\}$, and ${\mathsf{K}}_{R}$ communication receivers (simply called receivers), each receiver $k\in\mathcal{R}=\{1,\dots,{\mathsf{K}}_{R}\}$ with ${\mathsf{n}}_{k}$ antennas. The sensor is equipped with a single antenna. The channel outputs at the ${\mathsf{K}}_{R}$ receivers and at the sensor at the $t$-th time slot are,
@@
where ${\mathbf{H}}^{[ki]}(t)\in\mathbb{C}^{{\mathsf{n}}_{k}\times{\mathsf{m}}_{i}}$ represents the channel from the $i$-th transmitter to the $k$-th receiver at time slot $t$, and ${\mathbf{H}}^{[({\mathsf{K}}_{R}+1)i]}(t)\in\mathbb{C}^{1\times{\mathsf{m}}_{i}}$ represents the channel from the $i$-th transmitter to the sensor at time slot $t$; ${\mathbf{z}}^{[k]}(t)\in\mathbb{C}^{{\mathsf{n}}_{k}\times 1}$ and $z_{s}(t)\in\mathbb{C}$ are the additive Gaussian white noise at the receivers and sensor, respectively, with each element i.i.d. as $\mathcal{CN}(0,1)$. $g_{ki}\in\{0,1\}$ and $g_{s}\in\{0,1\}$ indicate the connectivity parameters between the transmitter and the receivers/sensor. The network connectivity (i.e., topology) is known to all the transmitters, receivers, and sensor. Assume that each transmission block is composed of $t_{0}$ channel uses. The signal transmitted by each transmitter $i\in\mathcal{T}$ at time slot $t\in\{1,2,\ldots,t_{0}\}$ is the sum of a communication signal and a dedicated sensing signal:
@@ II-A1 Heterogeneous coherence times
Consider that the transmitters and receivers/sensor are fully connected. The communication channel follows a block fading model with a short coherence time, whereas the sensing channel follows a block fading model with a long coherence time. The channel state information of the communication channel is perfectly known by the transmitters and receivers. [^6] Through channel estimation and prediction, that is, the trajectories of the receivers movements are regular and predictable. The difference in coherence time arises because the receivers (such as vehicular terminals) experience significant Doppler shifts due to high mobility, leading to a short channel coherence time (on the order of milliseconds). In contrast, sensor (such as static sensors) experiences negligible Doppler effects due to their low-speed or stationary nature, resulting in a long channel coherence time (tens of milliseconds or more), which supports continuous signal integration to enhance sensing accuracy. Under this scenario, we assume that the communication channel varies in every time slot following a Rayleigh fading model, while the sensor channel remains approximately constant over the given time period $t_{0}$.
@@
For the MU-MISO and MU-MIMO channels, an additional assumption is that the CSI of receivers are known to the transmitter in a non-causal manner [^7] There is a big body of literature in TDD/FDD massive MIMO systems[27, 28], where downlink channels can be efficiently estimated from uplink pilot signals by leveraging channel reciprocity in spatial and angular domains., implying that the transmitter has prior knowledge of the channel states for the entire block.
@@ II-A2 Heterogeneous connectivity
Now we consider the partial connectivity for network topology. Assume that the transmitters/sensor do not know the CSI of receivers and sensor, the receivers have perfect CSIR about their connected links. Both the receivers and sensor have channels with long coherence time, such that we assume that the channels remain constant during a single transmission block (i.e., $t_{0}$ time slots). Thus we can exploit the inherent properties of the network structure in a stable network to enhance the network performance.
We investigate two different scenarios, depending on whether the transmitters in the network can cooperate or not.
  • Non-cooperative transmitters. Each transmitter only has the communication messages desired by its corresponding receiver.
  • Cooperative transmitters. Each transmitter has all the communication messages desired by the receivers.
For the network topology, we consider two exiting partial connection networks, the regular network in [29, 30, 31, 26] and the neighboring antidotes network in [25, 32].
@@
The $({\mathsf{K}},d)$-regular network includes ${\mathsf{K}}$ single-antenna transmitters and ${\mathsf{K}}$ single-antenna users. Each user $j$ receives signals from transmitter $j$ as well as the successive $d-1$ ones, i.e., the set of connected transmitters by user $j$ is $\mathcal{R}_{j}=\{j,<j+1>_{{\mathsf{K}}},\dots,<j+d-1>_{{\mathsf{K}}}\}$. This model with regular linear connectivity (cells are enumerated sequentially) is also known as Wyner-type model for cellular networks and has been widely investigated in information theory[33, 34, 35].
The $({\mathsf{K}},U,D)$ neighboring antidotes network with ${\mathsf{K}}>U+D$ and $D\geq U$, includes ${\mathsf{K}}$ single-antenna transmitters and ${\mathsf{K}}$ single-antenna users. Each user $j$ is disconnected from the $U$ preceding transmitters (indexed $j-U,j-U+1,\dots,j-1$) and the $D$ succeeding transmitters (indexed $j+1,j+2,\dots,j+D$), where indices are interpreted modulo ${\mathsf{K}}$. User $j$ connects to all remaining ${\mathsf{K}}-U-D$ transmitters. Formally, $\mathcal{R}_{j}=\{j,<j+D+1>_{{\mathsf{K}}},\ldots,<j-U-1>_{{\mathsf{K}}}\}$. In addition, exchanging the values of $U$ and $D$ results in an equivalent network.
Based on the regular network in Definition 1, we consider two ways to add the sensor into the network:
  • Topology 1 (see Fig. 3(a)): In the $({\mathsf{K}}+1,d)$ regular network or the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network, we set the first ${\mathsf{K}}$ users as the communication receivers and the last user as the sensor.
  • Topology 2 (see Fig. 3(b)): In the $({\mathsf{K}},d)$ regular network or the $({\mathsf{K}},U,D)$ neighboring antidotes network, with the first ${\mathsf{K}}$ users being communication receivers and one additional user as the sensor, where the set of connected transmitters of the sensor is a subset of one communication communication receiver.
Define ${\mathcal{R}}^{[k]}_{c}$ and ${\mathcal{R}}_{s}$ as the set of connected transmitters by communication receiver $k$ and by the sensor, respectively.
Figure 3: Heterogeneous connectivity model: (a) Topology 1, (b) Topology 2.

### v5

@@ Abstract
Integrated sensing and communication (ISAC) systems provide significant enhancements in performance and resource efficiency compared to individual sensing and communication systems, primarily attributed to the collaborative use of wireless resources, radio waveforms, and hardware platforms.
This paper focuses on the bistatic ISAC systems with dispersed multi-receiver and one sensor. Compared to a monostatic ISAC system, the main challenge in the bistatic setting is that the information messages are unknown to the sensor and therefore they are seen as interference, while the channel between the transmitters (TX) and the sensor is unknown to the transmitters. In order to mitigate the interference at the sensor while maximizing the communication degree of freedom, we introduce two strategies, namely, blind interference alignment and topological interference management. Although well-known in the context of Gaussian interference channels, these strategies are novel in the context of bistatic ISAC.
For the bistatic ISAC models with heterogeneous coherence times or with heterogeneous connectivity, the achieved ISAC tradeoff points in terms of communication and sensing degrees of freedom are characterized. In particular, we show that the new tradeoff outperforms the time-sharing between the sensing-only and the communication-only schemes. Simulation results demonstrate that the proposed schemes significantly improve the channel estimation error for the sensing task, compared to treating interference as noise at the sensor and successive interference cancellation.
@@ I Introduction
Figure 1: The considered bistatic ISAC systems
@@
Despite the fundamental differences in information processing principles between sensing and communication, significant research efforts have been devoted to exploring the essence of “integration” gains in ISAC systems. These studies focus on the information-theoretic trade-offs between communication and sensing performance, aiming to establish a theoretical foundation for their co-design. The information-theoretic framework for ISAC was first introduced in [3, 4], where the authors investigated a monostatic ISAC system with state-dependent discrete memoryless channels and an independent and identically distributed (i.i.d.) state sequence. The capacity-distortion region was characterized by maximizing conditional mutual information under specific constraints. To explicitly derive the closed-form on the trade-off between communication and sensing performances, Xiong et al. [5] analyzed the Multiple-Input Multiple-Output (MIMO) Gaussian channel and explored the Cramér-Rao Bound (CRB)-rate region, which identified two extreme trade-off points: the communication-optimal point and the sensing-optimal point, which further revealed two critical aspects of ISAC trade-offs: subspace trade-offs and deterministic-random trade-offs. For scenarios with fixed sensing states, where the state remains constant over a transmission block, the optimal trade-off between communication rate and state detection error exponent was characterized in [6]. Monostatic ISAC over interference channels was considered in [7, 8], where two interfering ISAC transmitters communicate with their own users or a common user while sensing estimation through received echo signals, and an improved random coding than the superposition coding for interference channels was proposed. Readers can refer to the surveys in [9, 10] for more details on the information theoretic progress for ISAC.
@@
Most existing ISAC research has focused on monostatic systems, where sensing tasks exploit communication signals as side information. In addition, the monostatic architecture inherently suffers from strong self-interference and limited spatial coverage. In contrast, bistatic sensing—–where the sensor and transmitters are spatially separated—can avoid direct reflections, exploit multipath propagation, and enhance detection of low-altitude or ground targets [11]. From an information-theoretic perspective [^1] The authors are fully aware of the fact that a whole host of problems related to the synchronization (time, frequency, phase) of transmitters and sensors play an important role in bistatic sensing. Nevertheless, these problems can be circumvented by appropriate system design. For example, we may imagine that base stations and sensor are part of the same network running a high precision synchronization protocol, which effectively eliminate the problem to the required degree of accuracy. In contrast, conveying the information bits sent to the users by multiple base stations without a suitable delay is a much harder task, since these messages depend on local (at the base station) scheduling decisions and the throughput of such traffic would be huge. Hence, it is reasonable to assume that the information messages are not know at the sensor, while the synchronization issues have been reduced to a degree that they can be neglected for the sake of analytical tractability. , research on bistatic ISAC mainly concentrates on two scenarios depending on the relative positions of communication receivers(RX) and sensors.
@@
  • For the bistatic ISAC model with dispersed receivers and sensor, a key challenge arises from the fact that communication messages are not necessary to be decoded by the sensor and thus act as its interference. [^2] In this paper, we mainly focus on how to cancel the interference from the communication message to the sensor. Thus another existing bistatic ISAC model with dispersed receivers and sensor, where the sensor knows the communication message (such as in [16]), is out of scope. The authors in [17] developed three achievable decoding-and-estimation strategies for sensors, blind estimation (e.g., by treating interference as noise as in [18]), by partial-decoding-based estimation (e.g., by the successive interference cancellation (SIC) as in [19, 20]), and full-decoding-based estimation (i.e., fully decoding the communication message and then estimating). Interference mitigation via sharing radar sampling information and joint optimization of the transmit covariance matrices and radar sampling strategy was considered in [21], where the communication-induced interference was effectively mitigated to enable radar-communication coexistence. Mutual interference management through joint spatio-temporal transmit sequence design was studied in [22], where practical constraints such as power, Peak-to-Average Power Ratio (PAR), and spectral leakage were incorporated to ensure spectrum compatibility in MIMO coexistence systems. Physical secrecy issue was considered in [23], where the sensor cannot obtain any information about the communication message while estimating the state.
@@
This paper considers the bistatic ISAC model with multiple transmitters, dispersed multi-receivers, and one sensor. In order to characterize the fundamental tradeoff between communication and sensing performance, we consider “homogeneous” metrics, communication degree of freedom (cDoF) and sensing degree of freedom (sDoF), representing the average number of effective transmissions and observations in each time slot, respectively. Two new interference management strategies are first introduced into the bistatic ISAC model in order to perfectly eliminate interference to the sensor, instead of recovering all the interference as in [17] or reducing the interference as in [17, 21, 22].333 The interference alignment strategy was introduced into ISAC systems in [24], by using the original interference alignment scheme in [25] with the assumption that the channels from the communication transmitters and the (transformed) channels from the radar transmitters to the sensor are all known. In our paper, we assume that any channel to the sensor is unknown.
@@
  • One uses the blind interference alignment (BIA) strategy originally proposed in [26]. We assume that each transmitter is connected to all the receivers and sensor through wireless channels, and the channel between the transmitters and the sensor is slow-fading and unknown to the transmitters, while the channels between the transmitters and communication receivers are fast-fading and known to the transmitters and receivers.444 The practical motivation is twofold. (i) Communication channels vary rapidly due to user mobility, with a coherence time of approximately 1–10 ms. These channels are periodically estimated via UL/DL reciprocity and uplink pilots, from which the transmitters acquires channel state information (CSI). For simplicity, we assume perfect CSI and ignore pilot overhead. (ii) In contrast, sensing channels vary slowly as sensors and targets are typically stationary over short time spans. However, they are often passive and do not transmit pilots or provide feedback, making their CSI unavailable at the transmitters despite the slow variation. This asymmetric CSI assumption—known for communication but unknown for sensing—has also been adopted in information-theoretic ISAC models such as [5, 17]. We consider three types of wireless channels, including interference channel, multi-user multiple-input single-output (MU-MISO) channel, and multi-user MIMO (MU-MIMO) channel.
@@
  • The other strategy is the topological interference management (TIM) originally proposed in [27]. We follow the topological interference network model in [27, 28, 29], assuming that each transmitter is connected to a subset of the sensor and receivers through slow-fading wireless channels, while the transmitters only know the prior information of the network topology, instead of the full CSI. In this model, the regular network and the neighboring antidotes network are considered, determining the structure of the network topology.
@@
By using the two interference management strategies, BIA and TIM, we propose new tradeoff points (in closed-form) between cDoF and sDoF, which are strictly better than the time-sharing between the two extreme sensing-optimal and communication-optimal points. The common ingredient of using the above two strategies into ISAC is to leverage the channel heterogeneity among the sensor and receivers to align interference at the sensor without knowing the CSI of the sensor channel, where the heterogeneity is on the coherence times for the first model, and on the connection topology for the second model. Finally, we perform simulations on a practical ISAC system with Differential Quadrature Phase Shift Keying (DQPSK) modulation and additive white Gaussian noise. With channel estimation error as the sensing metric, the proposed method outperforms both treating interference as noise (TIN)-based sensing and SIC, which relies on decoding communication signals prior to sensing channel estimation.
The remainder of this paper is organized as follows. Section II introduces the system model of the considered ISAC system and the corresponding performance metrics. Section III summarizes the main results of this paper. Section IV describes the proposed ISAC schemes based on the BIA strategy, and Section V describes the proposed ISAC schemes based on the TIM strategy. Section VI provides simulation results to validate the performance of the proposed schemes. Finally, Section VII concludes the paper.
@@ II-A Channel Models
Consider a bistatic ISAC system, where ${\mathsf{K}}_{T}$ transmitters are equipped with ${\mathsf{m}}_{i}$ antennas where $i\in\mathcal{T}=\{1,\dots,{\mathsf{K}}_{T}\}$, and ${\mathsf{K}}_{R}$ communication receivers (simply called receivers), each receiver $k\in\mathcal{R}=\{1,\dots,{\mathsf{K}}_{R}\}$ with ${\mathsf{n}}_{k}$ antennas. Without loss of generality, assume that ${\mathsf{n}}_{1}\geq{\mathsf{n}}_{2}\geq\cdots\geq{\mathsf{n}}_{{\mathsf{K}}}$. The sensor is equipped with a single antenna. The channel outputs at the ${\mathsf{K}}_{R}$ receivers and at the sensor at the $t$-th time slot are,
@@
where ${\mathbf{H}}^{[ki]}(t)\in\mathbb{C}^{{\mathsf{n}}_{k}\times{\mathsf{m}}_{i}}$ represents the channel from the $i$-th transmitter to the $k$-th receiver at time slot $t$, and ${\mathbf{H}}^{[({\mathsf{K}}_{R}+1)i]}(t)\in\mathbb{C}^{1\times{\mathsf{m}}_{i}}$ represents the channel from the $i$-th transmitter to the sensor at time slot $t$; ${\mathbf{z}}^{[k]}(t)\in\mathbb{C}^{{\mathsf{n}}_{k}\times 1}$ and $z_{s}(t)\in\mathbb{C}$ are the additive Gaussian white noise at the receivers and sensor, respectively, with each element i.i.d. as $\mathcal{CN}(0,1)$. $g_{ki}\in\{0,1\}$ and $g_{s}\in\{0,1\}$ indicate the connectivity parameters between the transmitters and the receivers/sensor. The network connectivity (i.e., topology) is known to all the transmitters, receivers, and sensor. Assume that each transmission block is composed of $t_{0}$ channel uses. The signal transmitted by each transmitter $i\in\mathcal{T}$ at time slot $t\in\{1,2,\ldots,t_{0}\}$ is the sum of a communication signal and a dedicated sensing signal:
@@ II-A1 Heterogeneous coherence times
Consider that the transmitters and receivers/sensors are fully connected. The communication channel follows a block fading model with a short coherence time, whereas the sensing channel follows a block fading model with a long coherence time. The channel state information of the communication channel is perfectly known by the transmitters and receivers. [^6] Through channel estimation and prediction, that is, the trajectories of the receivers movements are regular and predictable. The difference in coherence time arises because the receivers (such as vehicular terminals) experience significant Doppler shifts due to high mobility, leading to a short channel coherence time (on the order of milliseconds). In contrast, sensors (such as static sensors) experience negligible Doppler effects due to their low-speed or stationary nature, resulting in a long channel coherence time (tens of milliseconds or more), which supports continuous signal integration to enhance sensing accuracy. Under this scenario, we assume that the communication channel varies in every time slot following a Rayleigh fading model, while the sensor channel remains approximately constant over the given time period $t_{0}$.
@@
For the MU-MISO and MU-MIMO channels, an additional assumption is that the CSI of receivers are known to the transmitters in a non-causal manner [^7] There is a big body of literature in TDD/FDD massive MIMO systems[30, 31], where downlink channels can be efficiently estimated from uplink pilot signals by leveraging channel reciprocity in spatial and angular domains., implying that the transmitters has prior knowledge of the entire block’s channel states.
@@ II-A2 Heterogeneous connectivity
Now we consider the partial connectivity for network topology. Assume that the receivers have perfect Channel State Information at the Receiver (CSIR) about their connected links and the sensor have no CSIR about the connected links.
Both the receivers and sensor have channels with long coherence time, such that we assume that the channels remain constant during a single transmission block (i.e., $t_{0}$ time slots). Thus we can exploit the inherent properties of the network structure in a stable network to enhance the network performance.
For the network topology, we consider two existing partial connection networks, the regular network in [32, 33, 34, 29] and the neighboring antidotes network in [28, 35]. These two topologies are widely adopted in topological interference management due to their structured yet flexible connectivity patterns. The $({\mathsf{K}},d)$-regular network, also known as the Wyner-type model, offers symmetric and localized interference, enabling tractable analysis and fundamental benchmarking of interference alignment schemes. On the other hand, the $({\mathsf{K}},U,D)$ neighboring antidotes network allows for asymmetric and generalized connectivity, effectively modeling heterogeneous interference scenarios where each receiver is protected from its closest interferers. Together, these models facilitate a comprehensive understanding of interference structures in ISAC systems under limited CSI assumptions.
The $({\mathsf{K}},U,D)$ neighboring antidotes network with ${\mathsf{K}}>U+D$ and $D\geq U$, includes ${\mathsf{K}}$ single-antenna transmitters and ${\mathsf{K}}$ single-antenna users. Each user $j$ is disconnected from the $U$ preceding transmitters (indexed $j-U,j-U+1,\dots,j-1$) and the $D$ succeeding transmitters (indexed $j+1,j+2,\dots,j+D$), where indices are interpreted modulo ${\mathsf{K}}$. User $j$ connects to all remaining ${\mathsf{K}}-U-D$ transmitters. Formally, $\mathcal{R}_{j}=\{j,<j+D+1>_{{\mathsf{K}}},\ldots,<j-U-1>_{{\mathsf{K}}}\}$. In addition, exchanging the values of $U$ and $D$ results in an equivalent network. For this topology, we consider non-cooperative transmitters, where each transmitter only has the communication messages desired by its corresponding receiver.
@@
The $({\mathsf{K}},d)$-regular network includes ${\mathsf{K}}$ single-antenna transmitters and ${\mathsf{K}}$ single-antenna users. Each user $j$ receives signals from transmitter $j$ as well as the successive $d-1$ ones, i.e., the set of connected transmitters by user $j$ is $\mathcal{R}_{j}=\{j,<j+1>_{{\mathsf{K}}},\dots,<j+d-1>_{{\mathsf{K}}}\}$. This model with regular linear connectivity (cells are enumerated sequentially) is also known as Wyner-type model for cellular networks and has been widely investigated in information theory [36, 37, 38]. For this topology, we consider cooperative transmitters, where each transmitter has all the communication messages desired by the receivers.
Based on the above networks, we consider the following way to add the sensor into the network: In the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network or the $({\mathsf{K}}+1,d)$ regular network, we set the first ${\mathsf{K}}$ users as the receivers and the last user as the sensor. Define ${\mathcal{R}}^{[k]}_{c}$ and ${\mathcal{R}}_{s}$ as the set of connected transmitters by communication receiver $k$ and by the sensor, respectively.
Figure 3: Heterogeneous connectivity model: (a) $({\mathsf{K}}+1,d)$-regular network, (b) $({\mathsf{K}}+1,U,D)$ neighboring antidotes network.